OpenStudy (anonymous):

satellite73 help me :) I got 643 as my answer for this question but I'm not sure if its right. I need someone to help double check. Question: Over the weekend, Trigg and Craig drove to Lake Pines to go fishing. Now they're preparing to go home. Trigg needs gas for his jeep, which gets 27 miles per gallon for gas millage. When he stops at the gas station he already has 5 gallons of gas in his tank. He buys more gas for $1.25 for galllon. How far will he go for $23.50 on gas. Round to nearest tenth. Please & thnk you :D

OpenStudy (anonymous):

guess you need to compute the amount of gas times gas mileage. i get \[(5+1.25\times 23.5)\times 27\]

OpenStudy (anonymous):

798.25?

OpenStudy (anonymous):

that is what i got yes!

OpenStudy (anonymous):

so rounded would be 798

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i got 928.125

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

/try what again rounding ?>

OpenStudy (anonymous):

did you compute \[(5+1.25\times 23.5)\times 27\]?

OpenStudy (anonymous):

no try the whole computation agani

OpenStudy (anonymous):

because i actually get 928.125 i tried it twice

OpenStudy (anonymous):

rounded to the nearest tenth would be 928.13

OpenStudy (anonymous):

are you sure ? i think you rounded to nearest 100th

OpenStudy (anonymous):

yes you are of course right. to the nearest tenth would be 928.1 i am lousy at rounding

OpenStudy (anonymous):

that was wrong :(

OpenStudy (anonymous):

really? let me try it again.

OpenStudy (anonymous):

its ok love :) so hi im jackie :)

OpenStudy (anonymous):

got the same answer again. hmmm

OpenStudy (anonymous):

do you know what the right answer is?

OpenStudy (anonymous):

hi jackie. why is this wrong any idea?

OpenStudy (anonymous):

i got 643 in the beginning

OpenStudy (anonymous):

i dont know really im going to try it again ; :) where are you from anyway ?

OpenStudy (anonymous):

are you from north carolina or new york o.O ; i tryed it again and it came out to my original anwser.

OpenStudy (radar):

Determine how many gallons of gas he bought with $23.50 when gas was $1.24 per gallon. (this was a long time ago lol)

OpenStudy (anonymous):

@radar says 1.25

OpenStudy (radar):

* $1.25 a gal.

OpenStudy (anonymous):

i did this like three times at least. still get the same answer. it is a trick or did i mess up somewhere?

OpenStudy (radar):

You multiplied when you should of divided first post.

OpenStudy (radar):

(23.50/1.25)+5 = amount of gas in gallons. that times 27 mpg will get it.

OpenStudy (anonymous):

heres another i could use help with ; the tools plus company is planning to produce drills. the company has rented a space for the manufacturing operation at $5000 per month. each drill requires $61 worth of materials and $36 worth of labor ; what is the final cost if the company maked 460 drills ? :)

OpenStudy (radar):

about 642.6 miles.

OpenStudy (anonymous):

god am i a moron.

OpenStudy (radar):

No, you're human.

OpenStudy (anonymous):

\[\color{red}{\text{ashamed}}\]

OpenStudy (anonymous):

@jackie last one drill requires 61+36 = $67 to make each one, plus 5000 so try \[97\times 460+5000\]

OpenStudy (anonymous):

aww :( dont be ashamed ! :) you only human love :) next time :) thank you anyway though :) i trust u :D

OpenStudy (anonymous):

45120 ?

OpenStudy (radar):

How many months are we talking about, is this one month production?

OpenStudy (anonymous):

dosent say /

OpenStudy (radar):

Lets assume and I think satellite did, it is only for one month.

OpenStudy (anonymous):

@satellite or 960$ & @radar it dosnet say moth really :( i wish it would ; make it alot easier .

OpenStudy (radar):

I think satellite got it, make the computation and that is probably the "approved answer"

OpenStudy (radar):

do the multiplication first, then the addition. Remember the rules for order of operation.

OpenStudy (anonymous):

so am i doing 61+36 = $67 to make each one, plus 5000 so try 97\times 460+5000

OpenStudy (radar):

I got $49,620 what you get

OpenStudy (anonymous):

45,120

OpenStudy (anonymous):

how did you get that ?

OpenStudy (radar):

97 times 460 =44,620 44,620 +5000=49620

OpenStudy (anonymous):

it wasnt correct :(

OpenStudy (radar):

Hmmm, have you stated the problem correctly? Is there something we are not considering?

OpenStudy (radar):

Did they only rent the building for one month?

OpenStudy (radar):

Sorry but I don't have a clue. Do you have what they want for an answer?